# Localization and splitting of a quantum droplet by immersing a heavy impurity

### v1

@@ III Static droplet configurations
First, we start by investigating the impact of the potential well of Eq. (2) on the harmonically trapped 1D droplet. Throughout this work, a weak harmonic trap of frequency $\Omega=0.001$ is used which naturally imposes an additional length scale, i.e. the harmonic oscillator length $l_{x}=\Omega^{-1/2}$ atop the free space droplet structure. This acts against the formation of spatially extended flat-top states Englezos et al. (2023) but also enforces an overall shift of the droplet’s binding energy towards less negative values Katsimiga et al. (2023a). Our aim, here, is to explore structural deformations, regions of existence and stability properties of the emergent droplets experiencing an additional central attractive or repulsive potential well which may emulate a heavy impurity atom inside the droplet. Below, our analysis for the stationary states mainly pertains to specific values of the chemical potential, $\mu=-0.2$, strength of the LHY attraction, $\delta=1$ and average mean-field repulsion $\lambda=1$. The fact that $\mu<0$ reflects the bound state character of the many-body state, while our focus is placed on $\left|V_{0}\right|\gg\left|\mu\right|$ such that the potential well significantly influences the droplet configuration. However, we have also verified the generalization of our results upon appropriate variations of these parameters (e.g. $\mu=-0.05,-0.1,-0.2222$, $\delta=0.7$) and suitable comments are provided within our description.
@@ III.1 Density distribution of the droplets
To visualize the structural modifications of the harmonically trapped 1D droplet due to the presence of the potential well we inspect the density, $n(x)=\left|\psi(x)\right|^{2}$, of the system for different barrier characteristics, see Fig. 2. As discussed in Sec. II and depicted in Fig. 1(f) the attractive potential well accommodates additional bound states whose number increases for larger $V_{0}<0$. Hence, atoms from the droplet accumulate in these bound states substantially distorting its spatial distribution. Prototype examples confirming this behavior are shown in Figs. 2(a)-(c) for $\mu=-0.2$ and distinct values of $\sigma$ in terms of $V_{0}$. Specifically, the original (at $V_{0}=0$) flat-top droplet gradually shrinks in size and its peak density decreases for more negative values of $V_{0}$ since a more appreciable amount of atoms are trapped within the central potential dip. In fact, the computed structurally deformed droplet configurations follow the analytical solution reported in Refs. Debnath et al. (2023); Abdullaev and Galimzyanov (2020) in the presence of a delta potential
@@
where $\xi_{D}$ is the amended droplet’s healing length. We have confirmed this excellent agreement (not shown) upon fitting Eq. (4) to the numerically obtained non-zero eGPE waveforms presented in Fig. 2(a)-(c). The above-mentioned trapping property depends crucially on the height of the central attractive potential as the number of bound states at the single-particle level increases for larger $V_{0}$ and fixed $\sigma$, $\mu$, see also Fig. 1(f). Hence, the effect of the central attractive potential is more prominent for larger $\sigma$, compare in particular Figs. 2(a), (b) and (c), leading eventually to the termination of the droplet solution. For instance, in the case of $\mu=-0.2$, $\delta=1$, displayed in Fig. 2(a)-(c), the droplet disappears at $\left|V_{0}\right|=5,~{}2.7,~{}0.68$ when $\sigma=0.05,~{}0.1,~{}0.5$ respectively. The above thresholds match closely the ones derived for a droplet in a Dirac delta potential, which exists only for $-\mu>\left|V_{0}\right|^{2}\sigma^{2}\pi$ Debnath et al. (2023).
@@
A similar response takes also place for different values of the chemical potential (not shown) and the LHY strength as long as the droplet solution exists. Recall that the latter occurs only for $\left|\mu\right|>2\delta^{2}/9$ Katsimiga et al. (2023a) and thus for $\delta=0.7$ for instance, droplets exist at $\mu=-0.1$. For these parameters, the impact of the attractive potential on the droplet [inset of Fig. 2(c)] is less drastic compared to $\delta=1$, leading to the destruction of the structure for relatively larger values of $V_{0}$. For example, fixing $\mu=-0.1$ we find that the solution terminates at $V_{0}\approx-0.42$ [$V_{0}\approx-1.78$] for $\delta=0.7$ and at $V_{0}=-0.40$ [$V_{0}=-1.74$] for $\delta=1$ when $\sigma=0.05\ll\xi$ [$\sigma=0.1<\xi$]. This is because the droplet distribution for a specific $\mu$ becomes broader and less peaked as $\delta$ decreases due to the smaller 1D LHY attraction (with the mean-field repulsion dominating). Thus, it requires a larger attractive $V_{0}$ to achieve the same number of atoms in the bound states for broader droplet distributions.
@@
On the other hand, the free droplet solution in the absence of the central potential has been demonstrated to exist for $-0.222<\mu<0$ Astrakharchik and Malomed (2018); Katsimiga et al. (2023a) when $\delta=1$. In this case, keeping all parameters fixed and tuning $\mu$ towards the lower bound a transition from Gaussian type to flat-top droplet configurations occurs. As such droplets characterized by less negative $\mu$ are more susceptible to the presence of the additional potential dip and cease to exist for smaller negative values of $\mu$ since they accommodate a relatively smaller amount of atoms as was shown, for instance, in Refs. Astrakharchik and Malomed (2018); Katsimiga et al. (2023a).
@@
Figure 3: (a) Real part of the BdG excitation spectra for the harmonically trapped droplet under the influence of an external attractive potential well as a function of its height, $V_{0}$. The behavior of the collective excitation branches, labeled $\omega^{(n)}_{R}$ with $n=0,1,2,\ldots$, is shown. Selected imaginary versus real parts of the ensuing BdG droplet spectra in the case of (b)-(d) an attractive and (e)-(g) a repulsive potential for different potential characteristics (see legends). The presence (absence) of finite imaginary parts in the case of a repulsive (an attractive) potential reflects the unstable (stable) nature of the droplet configurations. (h), (i) Dynamical evolution of the perturbed stationary configurations at selected $V_{0}$ values of the repulsive potential. In all cases, the chemical potential, $\mu=-0.2$, and the LHY strength is $\delta=1$ (see legends).
@@
Turning to repulsive central potentials, characterized by $V_{0}>0$, we observe a completely different effect on the droplet distribution. Ramping-up the repulsive barrier progressively enforces a splitting of the single droplet. Initially, this process yields a shallow two-hump configuration [see, e.g., Fig. 2(d) for $40.5<V_{0}<$ 4] leading eventually to a well-separated two droplet distribution [see, for instance, Fig. 2(d) for $V_{0}>5$]. As can be seen from the density distributions in Figs. 2(d)-(f), the aforementioned splitting mechanism occurs naturally earlier in terms of $V_{0}$ for larger widths because the exerted force from the barrier to the droplet is increased.
@@
$\Psi_{{\rm D2}}(x,\mu)=\frac{3\mu}{\delta\big{(}1+\sqrt{1+\frac{9\mu}{2\delta^{2}}}\cosh(\sqrt{-2\mu}(x-x_{0})\big{)}}.$ (5)
Here, $\mu$ denotes the background chemical potential, $x_{0}$ is the center of each droplet segment and $\delta$ refers to the LHY strength. These parameters are determined by the fitting to the eGPE waveform. An excellent agreement between our numerically obtained two-droplet solution and the analytical estimate can be extracted (not shown for brevity), thus confirming the droplet character of the former. A careful inspection of Figs. 2(d)-(f) also reveals that the participating density fragments in the two-droplet configurations are thicker for reduced potential width, compare for instance Figs. 2(d) and (f). This is expected since in the case of a “thicker” barrier the local harmonic oscillators possess tighter frequencies. As such, the trapping becomes more prominent resulting in relatively shrinked droplets.
@@ III.2 Stability analysis of the droplet structures
Having analyzed the deformations of the droplet configurations due to the presence of the central potential well and their parametric regions of existence, we next examine the spectral stability of the ensuing solutions through a BdG analysis Stoof et al. (2009). Note here that the excitation spectrum of symmetric droplets has been previously discussed in 1D Tylutki et al. (2020); Katsimiga et al. (2023b) and 2D Fei et al. (2024); Bougas et al. (2024) but also for genuine two-component droplets Charalampidis and Mistakidis (2025) in both free space or under the influence of a harmonic trap, see also Refs. Katsimiga et al. (2023a); Liu et al. (2022) for estimating the impact of nonlinear excitations. On the other hand, the stability of droplets experiencing a potential well atop the harmonic trap remains unprecedented to the best of our knowledge.
@@
In this expression, [$a(x)$, $b(x)$] represents the eigenvector with eigenfrequency $\omega=\omega_{R}+i\,\omega_{I}$. Inserting the aforementioned ansatz into Eq. (3) and keeping terms of $\mathcal{O}(\varepsilon)$ allows the extraction of a linearized set of equations that has been reported, for instance, in Refs. Tylutki et al. (2020); Katsimiga et al. (2023a, b). The relevant equations are subsequently solved numerically yielding the ensuing discrete eigenfrequencies, due to the presence of the external trap, and eigenvectors providing access to the spectrum of the system and thus allowing us to deduce its stability properties. It is apparent that a configuration is deemed stable as long as $\omega_{I}=0$; otherwise instabilities occur. On the other hand, $\omega_{R}$ captures the collective excitation frequencies of the respective configuration and their behavior as a function of a system parameter.
Figure 4: Density evolution of the droplet after ramping-up the central repulsive potential of height $V_{0}=1$ and width (a) $\sigma=0.05$, (d) $\sigma=0.5$. (b) [(e)] Selected density profiles taken from (a) [(d)] at specific time-instants (see legends), along with the droplet fits (dashes lines) obtained through Eq. (5). The quench tends to produce two counterpropagating droplet fragments. The latter move further apart and their separation $D(t)$ increases both by (c) varying $V_{0}$ and fixing the width (see legend), and (f) varying $\sigma$ while $V_{0}=1$. All other parameters are the same as in Fig. 3.
@@
The real part of the BdG spectrum, $\omega_{R}$, is illustrated in Fig. 3(a) for droplet distributions with $\mu=-0.2$ and $\delta=1$ upon parametric variation of the height of the attractive potential well characterized by a width $\sigma=0.05\ll\xi$. The dependence of the first few eigenfrequencies, $\omega^{(n)}_{R}$ with $n=0,1,\ldots,6$, on the potential depth (or height) can be discerned. Specifically, the trajectory traced by $\omega^{(0)}_{R}=0$ corresponds to the zero-eigenfrequency mode that is related to an infinitesimal phase shift of the droplet’s wave function (phase invariance). The consecutive one, i.e. $\omega^{(1)}_{R}$, is associated with the droplet’s center-of-mass displacement (dipole mode Tylutki et al. (2020); Englezos et al. (2023)) and the trajectory traced by $\omega^{(2)}_{R}$ corresponds to the collective breathing mode Tylutki et al. (2020); Astrakharchik and Malomed (2018) of the droplet solutions. For increasing $V_{0}$ while keeping $\sigma$ fixed, it turns out that $\omega^{(n)}_{R}$ are shifted to lager values. For example, $\omega^{(2)}_{R}\approx 0.181$ for $V_{0}=-1$ becoming $\omega^{(2)}_{R}\approx 0.214$ for $V_{0}=-4.99$ [Fig.3 (a)] at the termination point of the droplet solution. The same holds true upon fixing $V_{0}$ e.g. to $V_{0}=-0.2$ and varying $\sigma$. In this case, we find that $\omega^{(2)}_{R}\approx 0.155$ for $\sigma=0.05$, acquiring the value of $\omega^{(2)}_{R}\approx 0.197$ for $\sigma=0.5$.
@@
Importantly, all solutions identified herein stemming from an attractive central potential are stable irrespective of its width and height. This is verified by the zero imaginary part of the eigenfrequencies. To explicate this general property we present selective eigenspectra in the $\omega_{R}-\omega_{I}$ plane for different barrier heights and widths in Figs. 3(b)-(d). It can be readily seen that in all cases $\omega_{I}=0$ (being typically of the order of $\approx 10^{-8}$ herein) which supports the spectral stability of these deformed droplet solutions. To independently verify the stability predicted by the BdG analysis, we confirmed that the solutions remain intact when freely evolved up to $t=10^{4}$ (in dimensionless units).
@@
Turning to the repulsive potential case, a completely different deformation of the droplet configuration occurs. Recall that as we depart from $V_{0}=0$, namely from the harmonically trapped state, the original droplet gets depleted, with a gradually increasing density dip being imprinted around its core due to the finite and progressively higher value of $V_{0}$. This depletion leads to the segregation of the stationary configuration into two fragments whose separation increases for larger $V_{0}$ [see Fig.2(d)-(f)]. This deformed configuration, is found to be unstable for all values of $V_{0}$ [Fig. 3(e)-(g)]. The maximum growth rate of the ensuing instability corresponds to $\omega_{I}=0.044$ occurring at $V_{0}=0.124$ e.g. for $\sigma=0.5$. The relevant real vs imaginary BdG spectrum is presented in Fig. 3(f) where the imaginary eigenfrequency pair can be seen. However, as segregation progresses, the relevant eigenfrequencies acquire $\omega_{I}\sim\mathcal{O}(10^{-4})$ [Fig. 3(g)] asymptotically tending to two fairly separated stable individual droplets (see also the discussion below).
@@
Figure 5: Dynamics of the droplet density following a quench of the central attractive potential from height $V_{0}=0$ to $V_{0}=-1$ for various widths (a) $\sigma=0.05$ and (c) $\sigma=0.5$. Relative small amplitude quenches trigger a collective motion of the droplet [panel (a)]. However, larger amplitudes lead to self-evaporation along with droplet excitation at short times which eventually decays tending to a quasi-steady state at long times [panel (c)]. (b) [(d)] Instantaneous densities taken from (a) [(c)]. Panels (b), (d) demonstrate also fits of the analytical droplet solution described by Eq. (4) and Eq. (5) respectively, to the eGPE obtained wave functions at specific time-instants (see legend). Adequate agreement occurs, excluding the tails of the distributions. The chemical potential used is $\mu=-0,2$ and the strength of the LHY $\delta=1$.

### v2

@@ III Static droplet configurations
First, we start by investigating the impact of the potential well of Eq. (2) on the harmonically trapped 1D droplet. Throughout this work, a weak harmonic trap of frequency $\Omega=0.001$ is used which naturally imposes an additional length scale, i.e. the harmonic oscillator length $l_{x}=\Omega^{-1/2}$ atop the free space droplet structure. This acts against the formation of spatially extended flat-top states Englezos et al. (2023) but also enforces an overall shift of the droplet’s binding energy towards less negative values Katsimiga et al. (2023a). Our aim, here, is to explore structural deformations, regions of existence and stability properties of the emergent droplets experiencing an additional central potential well or barrier which may emulate a heavy impurity atom inside the droplet. Below, our analysis for the stationary states mainly pertains to specific values of the chemical potential, $\mu=-0.2$, strength of the LHY attraction, $\delta=1$ and average mean-field repulsion $\lambda=1$. The fact that $\mu<0$ reflects the bound state character of the many-body state, while our focus is placed on $\left|V_{0}\right|\gg\left|\mu\right|$ such that the potential defect significantly influences the droplet configuration. However, we have also verified the generalization of our results upon appropriate variations of these parameters (e.g. $\mu=-0.05,-0.1,-0.2222$, $\delta=0.7$) and suitable comments are provided within our description.
@@ III.1 Density distribution of the droplets
To visualize the structural modifications of the harmonically trapped 1D droplet due to the presence of the potential well we inspect the density, $n(x)=\left|\psi(x)\right|^{2}$, of the system for different barrier characteristics, see Fig. 2. As discussed in Sec. II and depicted in Fig. 1(f) the potential well accommodates additional bound states whose number increases for larger $V_{0}<0$. Hence, atoms from the droplet accumulate in these bound states substantially distorting its spatial distribution. Prototype examples confirming this behavior are shown in Figs. 2(a)-(c) for $\mu=-0.2$ and distinct values of $\sigma$ in terms of $V_{0}$. Specifically, the original (at $V_{0}=0$) flat-top droplet gradually shrinks in size and its peak density decreases for more negative values of $V_{0}$ since a more appreciable amount of atoms are trapped within the central potential well. In fact, the computed structurally deformed droplet configurations follow the analytical solution reported in Refs. Debnath et al. (2023); Abdullaev and Galimzyanov (2020) in the presence of a delta potential
@@
where $\xi_{D}$ is the amended droplet’s healing length. We have confirmed this excellent agreement (not shown) upon fitting Eq. (4) to the numerically obtained non-zero eGPE waveforms presented in Fig. 2(a)-(c). The above-mentioned trapping property depends crucially on the height of the central potential well as the number of bound states at the single-particle level increases for larger $V_{0}$ and fixed $\sigma$, $\mu$, see also Fig. 1(f). Hence, the effect of the central potential well is more prominent for larger $\sigma$, compare in particular Figs. 2(a), (b) and (c), leading eventually to the termination of the droplet solution. For instance, in the case of $\mu=-0.2$, $\delta=1$, displayed in Fig. 2(a)-(c), the droplet disappears at $\left|V_{0}\right|=5,~2.7,~0.68$ when $\sigma=0.05,~0.1,~0.5$ respectively. The above thresholds match closely the ones derived for a droplet in a Dirac delta potential, which exists only for $\mu>\left|V_{0}\right|^{2}\sigma^{2}\pi$ Debnath et al. (2023).
@@
A similar response takes also place for different values of the chemical potential (not shown) and the LHY strength as long as the droplet solution exists. Recall that the latter occurs only for $\left|\mu\right|>2\delta^{2}/9$ Katsimiga et al. (2023a) and thus for $\delta=0.7$ for instance, droplets exist at $\mu=-0.1$. For these parameters, the impact of the potential well on the droplet [inset of Fig. 2(c)] is less drastic compared to $\delta=1$, leading to the destruction of the structure for relatively larger values of $V_{0}$. For example, fixing $\mu=-0.1$ we find that the solution terminates at $V_{0}\approx-0.42$ $V_{0}\approx-1.78$] for $\delta=0.7$ and at $V_{0}=-0.40$ $V_{0}=-1.74$] for $\delta=1$ when $\sigma=0.05\ll\xi$ $\sigma=0.1<\xi$]. This is because the droplet distribution for a specific $\mu$ becomes broader and less peaked as $\delta$ decreases due to the smaller 1D LHY attraction (with the mean-field repulsion dominating). Thus, it requires a larger negative $V_{0}$ to achieve the same number of atoms in the bound states for broader droplet distributions.
@@
On the other hand, the free droplet solution in the absence of the central potential has been demonstrated to exist for $0.222<\mu<0$ Astrakharchik and Malomed (2018); Katsimiga et al. (2023a) when $\delta=1$. In this case, keeping all parameters fixed and tuning $\mu$ towards the lower bound a transition from Gaussian type to flat-top droplet configurations occurs. As such droplets characterized by less negative $\mu$ are more susceptible to the presence of the additional potential defect and cease to exist for smaller negative values of $\mu$ since they accommodate a relatively smaller amount of atoms as was shown, for instance, in Refs. Astrakharchik and Malomed (2018); Katsimiga et al. (2023a).
@@
Figure 3: (a) Real part of the BdG excitation spectra for the harmonically trapped droplet under the influence of an external potential well as a function of its height, $V_{0}$. The behavior of the collective excitation branches, labeled $\omega^{(n)}_{R}$ with $n=0,1,2,\ldots$, is shown. Selected imaginary versus real parts of the ensuing BdG droplet spectra in the case of (b)-(d) potential well and (e)-(g) a potential barrier for different potential characteristics (see legends). The presence (absence) of finite imaginary parts in the case of a potential barrier (well) reflects the unstable (stable) nature of the droplet configurations. (h), (i) Dynamical evolution of the perturbed stationary configurations at selected $V_{0}$ values of the potential barrier. In all cases, the chemical potential, $\mu=-0.2$, and the LHY strength is $\delta=1$ (see legends).
@@
Turning to central potential barriers, characterized by $V_{0}>0$, we observe a completely different effect on the droplet distribution. Ramping-up the barrier progressively enforces a splitting of the single droplet. Initially, this process yields a shallow two-hump configuration [see, e.g., Fig. 2(d) for $0.5<V_{0}<4$] leading eventually to a well-separated two droplet distribution [see, for instance, Fig. 2(d) for $V_{0}>5$]. As can be seen from the density distributions in Figs. 2(d)-(f), the aforementioned splitting mechanism occurs naturally earlier in terms of $V_{0}$ for larger widths because the exerted force from the barrier to the droplet is increased.
@@
$\Psi_{{\rm D2}}(x,\mu)=\frac{3\mu}{\delta\left(1+\sqrt{1+\frac{9\mu}{2\delta^{2}}}\cosh\left(\sqrt{-2\mu}(x-x_{0})\right)\right)}.$ (5)
Here, $\mu$ denotes the background chemical potential, $x_{0}$ is the center of each droplet segment and $\delta$ refers to the LHY strength. These parameters are determined by the fitting to the eGPE waveform. An excellent agreement between our numerically obtained two-droplet solution and the analytical estimate can be extracted (not shown for brevity), thus confirming the droplet character of the former. A careful inspection of Figs. 2(d)-(f) also reveals that the participating density fragments in the two-droplet configurations are thicker for reduced potential barrier width, compare for instance Figs. 2(d) and (f). This is expected since in the case of a “thicker” barrier the local harmonic oscillators possess tighter frequencies. As such, the trapping becomes more prominent resulting in relatively shrinked droplets.
@@ III.2 Stability analysis of the droplet structures
Having analyzed the deformations of the droplet configurations due to the presence of the central potential defect and their parametric regions of existence, we next examine the spectral stability of the ensuing solutions through a BdG analysis Stoof et al. (2009). Note here that the excitation spectrum of symmetric droplets has been previously discussed in 1D Tylutki et al. (2020); Katsimiga et al. (2023b) and 2D Fei et al. (2024); Bougas et al. (2024) but also for genuine two-component droplets Charalampidis and Mistakidis (2025) in both free space or under the influence of a harmonic trap, see also Refs. Katsimiga et al. (2023a); Liu et al. (2022) for estimating the impact of nonlinear excitations. On the other hand, the stability of droplets experiencing a central potential defect atop the harmonic trap remains unprecedented to the best of our knowledge.
@@
In this expression, [$a(x)$, $b(x)$] represents the eigenvector with eigenfrequency $\omega=\omega_{R}+i\,\omega_{I}$. Here, $\omega_{R}$ ($\omega_{I}$) corresponds to the real (imaginary) part of the eigenfrequency. Inserting the aforementioned ansatz into Eq. (3) and keeping terms of $\mathcal{O}(\varepsilon)$ allows the extraction of a linearized set of equations that has been reported, for instance, in Refs. Tylutki et al. (2020); Katsimiga et al. (2023a, b). The relevant equations are subsequently solved numerically yielding the ensuing discrete eigenfrequencies, due to the presence of the external trap, and eigenvectors providing access to the spectrum of the system and thus allowing us to deduce its stability properties. It is apparent that a configuration is deemed stable as long as $\omega_{I}=0$; otherwise instabilities occur. On the other hand, $\omega_{R}$ captures the collective excitation frequencies of the respective configuration and their behavior as a function of a system parameter.
@@
The real part of the BdG spectrum, $\omega_{R}$, is illustrated in Fig. 3(a) for droplet distributions with $\mu=-0.2$ and $\delta=1$ upon parametric variation of the height of the potential well characterized by a width $\sigma=0.05\ll\xi$. The dependence of the first few eigenfrequencies, $\omega^{(n)}_{R}$ with $n=0,1,\ldots,6$, on the potential depth (or height) can be discerned. Specifically, the trajectory traced by $\omega^{(0)}_{R}=0$ corresponds to the zero-eigenfrequency mode that is related to an infinitesimal phase shift of the droplet’s wave function (phase invariance). The consecutive one, i.e. $\omega^{(1)}_{R}$, is associated with the droplet’s center-of-mass displacement (dipole mode Tylutki et al. (2020); Englezos et al. (2023)) and the trajectory traced by $\omega^{(2)}_{R}$ corresponds to the collective breathing mode Tylutki et al. (2020); Astrakharchik and Malomed (2018) of the droplet solutions. For increasing $V_{0}$ while keeping $\sigma$ fixed, it turns out that $\omega^{(n)}_{R}$ are shifted to larger values. For example, $\omega^{(2)}_{R}\approx 0.181$ for $V_{0}=-1$ becoming $\omega^{(2)}_{R}\approx 0.214$ for $V_{0}=-4.99$ [Fig.3 (a)] at the termination point of the droplet solution. The same holds true upon fixing $V_{0}$ e.g. to $V_{0}=-0.2$ and varying $\sigma$. In this case, we find that $\omega^{(2)}_{R}\approx 0.155$ for $\sigma=0.05$, acquiring the value of $\omega^{(2)}_{R}\approx 0.197$ for $\sigma=0.5$.
@@
Importantly, all solutions identified herein stemming from a potential well are stable irrespective of its width and height. This is verified by the zero imaginary part of the eigenfrequencies. To explicate this general property we present selective eigenspectra in the $\omega_{R}-\omega_{I}$ plane for different potential well heights and widths in Figs. 3(b)-(d). It can be readily seen that in all cases $\omega_{I}=0$ (being typically of the order of $\approx 10^{-8}$ herein) which supports the spectral stability of these deformed droplet solutions. To independently verify the stability predicted by the BdG analysis, we confirmed that the solutions remain intact when freely evolved up to $t=10^{4}$ (in dimensionless units).
@@
Turning to the potential barrier case, a completely different deformation of the droplet configuration occurs. Recall that as we depart from $V_{0}=0$, namely from the harmonically trapped state, the original droplet gets depleted, with a gradually increasing density dip being imprinted around its core due to the finite and progressively higher value of $V_{0}$. This depletion leads to the segregation of the stationary configuration into two fragments whose separation increases for larger $V_{0}$ [see Fig.2(d)-(f)]. This deformed configuration, is found to be unstable for all considered non-negligible values of $V_{0}$ [see e.g. Fig. 3(e)-(g)] which alter the original droplet structure. The maximum growth rate of the ensuing instability corresponds to $\omega_{I}=0.044$ occurring at $V_{0}=0.124$ e.g. for $\sigma=0.5$. The relevant real vs imaginary BdG spectrum is presented in Fig. 3(f) where the imaginary eigenfrequency pair can be seen. However, as segregation progresses, the relevant eigenfrequencies acquire $\omega_{I}\sim\mathcal{O}(10^{-4})$ [Fig. 3(g)] asymptotically tending to two fairly separated stable individual droplets (see also the discussion below).
@@
Figure 4: Density evolution of the droplet after ramping-up the central potential barrier of height $V_{0}=1$ and width (a) $\sigma=0.05$, (d) $\sigma=0.5$. (b) [(e)] Selected density profiles taken from (a) [(d)] at specific time-instants (see legends), along with the droplet fits (dashes lines) obtained through Eq. (5). The quench tends to produce two counterpropagating droplet fragments. The latter move further apart and their separation $D(t)$ increases both by (c) varying $V_{0}$ and fixing the width (see legend), and (f) varying $\sigma$ while $V_{0}=1$. All other parameters are the same as in Fig. 3.